Over a season in a​ women's basketball league Jackson scored 29 more points than the​ second-highest scorer, Leslie.​ Together, Jackson and Leslie scored 1139 points during the season. How many points did each player score over the course of the​ season?
Jackson scored points?
Leslie scored points?

Answer by stanbon(75887)

Over a season in a​ women's basketball league Jackson scored 29 more points than the​ second-highest scorer, Leslie.​ Together, Jackson and Leslie scored 1139 points during the season. How many points did each player score over the course of the​ season?
Equations::
J - L = 29
J + L = 1139
------
Add and solve for "J":
2J = 1168
J = 584 (Jackson points)
L = 1139 - 584 = 555 (Leslie points)
--------------
Cheers,
Stan H.
-------------